Circular No. 6996

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1998 Q1 (LINEAR) Preliminary parabolic orbital elements by the undersigned from 23

observations, Aug. 24–25:

T = 1998 June 28.131 TT  ω = 133.^{o}338
Ω = 159.831 }2000.0
q = 1.56032 AU  i = 32.021
1998 TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Aug. 15 22^{h}05.^{m}22 + 7^{o}34.^{′}7 0.712 1.692 157.^{o}2 13.^{o}4 15.5
20 22 07.35 + 3 39.1 0.725 1.719 163.5  9.6 15.7
25 22 09.29 − 0 13.6 0.745 1.748 169.3  6.2 15.8
30 22 11.14 − 3 56.3 0.773 1.779 172.5  4.2 15.9
Sept. 4 22 13.03 − 7 23.5 0.810 1.812 170.6  5.2 16.1
9 22 15.04 −10 31.1 0.854 1.847 165.7  7.7 16.3
14 22 17.24 −13 17.0 0.905 1.883 160.2 10.4 16.5
19 22 19.67 −15 40.5 0.963 1.920 154.7 12.9 16.8
24 22 22.40 −17 42.1 1.027 1.959 149.4 15.1 17.0

SUPERNOVAE 1998do AND 1998dp L. Germany reports that the Mount Stromlo Abell Cluster Supernova

Search Team (cf. IAUC 6639) has discovered two apparent supernovae on V and R CCD images taken by A. Graham with the Mount Stromlo 1.27-m telescope (+ Macho Camera), as follows:

SN  1998 UT  α_{2000}  δ_{2000}  V  Offset
1998do Aug. 1 1^{h}14^{m}30.^{s}92  + 0^{o}28′13′′.0 19.0 0′′.2 E, 0′′.4 N
1998dp Aug. 14 21 58 28.12  −19 58 05.9 20.3  0′′.6 N

The limiting V magnitudes for the observations were V ∼ 19.5 and 20.5 on Aug. 1 and 14, respectively. Confirmation images were taken by A. Graham with the same telescope on Aug. 22 (limiting mag V ∼ 20.5). SNe 1998do and 1998dp were not visible on images taken on 1997 Dec. 4 (limiting mag V ∼ 19.5) and on 1998 June 8 (limiting mag V ∼ 20.5), respectively. SN 1998do is near Abell 168, and SN 1998dp is near Abell 2401.

1998 August 25  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green